# Revisiting boundary-driven method for transport: Finite-size effects and the role of system-bath coupling

### v1

@@ Abstract
Understanding transport in interacting quantum many-body systems is a central challenge in condensed matter and statistical physics. Numerical studies typically rely on two main approaches: Dynamics of linear-response functions in closed systems and Markovian dynamics governed by master equations for boundary-driven open systems. While the equivalence of their dynamical behavior has been explored in recent studies, a systematic comparison of the transport coefficients obtained from these two classes of methods remains an open question. Here, we address this gap by comparing and contrasting the dc diffusion constant $\mathcal{D}_{\text{dc}}$ computed from the aforementioned two approaches. We find a clear mismatch between the two, with $\mathcal{D}_{\text{dc}}$ exhibiting a strong dependence on the system-bath coupling for the boundary-driven technique, highlighting fundamental limitations of such a method in calculating the transport coefficients related to asymptotic dynamical behavior of the system. We trace the origin of this mismatch to the incorrect order of limits of time $t\rightarrow\infty$ and system size $L\rightarrow\infty$, which we argue to be intrinsic to boundary-driven setups. As a practical resolution, we advocate computing only time-dependent transport coefficients within the boundary-driven framework, which show excellent agreement with those obtained from the Kubo formalism based on closed-system dynamics, up to a timescale set by the system size. This leads us to interpret the sensitivity of the dc diffusion constant on the system-bath coupling strength in an open system as a potential diagnostic for finite-size effects.
@@ I Introduction
Over the last few decades, studying nonequilibrium dynamics in quantum many-body systems has become a key area of research in both theoretical and experimental condensed matter physics, with broad ramifications ranging from fundamental questions in statistical mechanics to advanced applications in quantum control and quantum computation [1, 2, 3, 4]. One of the most ubiquitous manifestations of nonequilibrium behavior are transport processes of conserved quantities such as energy, particle number and magnetization, which are relevant for both closed/isolated and open quantum systems coupled to external environments. A comprehensive understanding of transport is thus crucial for elucidating on the one hand the emergent macroscopic behavior as well as on the other hand the underlying microscopic mechanisms governing quantum many-body dynamics such as the build up of nonequilibrium steady states (NESS) and the relaxation to them [5]. Over the past years, notable progress in this field has been driven by the development of fresh theoretical concepts like eigenstate thermalization hypotheses [6, 7, 8, 9] and dynamical quantum typicality [10, 11, 12, 13, 14], as well as sophisticated analytical and numerical methods and advanced experimental techniques. We now provide a brief overview of some of the key numerical and analytical methods commonly employed to study such problems, which are particularly relevant to the present work.
@@
On the numerical frontier, another alternative to traditional LRT-based closed-system technique for studying transport is given within the framework of boundary-driven open quantum systems. In such a setup, the system is coupled to an environment, for instance, modeled by external baths at different temperatures or chemical potentials [26, 27, 28, 29, 30]. This coupling drives the system out of equilibrium and induces transport, typically resulting in a NESS with a constant current in the bulk and a characteristic density profile in the long-time limit. A widely employed approach for describing the dynamics of such open-system scenarios is provided by Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) quantum master equation [31, 32]. While the derivation of the GKSL master equation from a microscopic system-bath model is a nontrivial task in practice [33, 34], it provides the most general form of a time-local (Markovian) master equation whose dynamical map is completely positive and trace-preserving (CPTP), thus ensuring the evolution of any physical density matrix into another valid density matrix [35]. This formalism offers considerable advantages for numerical simulations, particularly in conjunction with stochastic methods or matrix-product-state techniques based upon vectorized density matrices.
@@
Using advanced numerical techniques such as stochastic unraveling (SU) and time-evolving block decimation (TEBD), we uncover a significant discrepancy in the long-time dc diffusion constant $\mathcal{D}_{\text{dc}}$ obtained from the Kubo formalism (closed system) versus boundary-driven open systems in finite systems. In particular, $\mathcal{D}_{\text{dc}}$ extracted from the NESS in boundary-driven setups exhibits a strong dependence on the system-bath coupling strength $\gamma$, revealing a fundamental limitation of this method in capturing asymptotic transport properties. To pinpoint the origin of this issue, we compute the time-dependent diffusion constant $\mathcal{D}(t)$, which agrees rather well with Kubo-based results from closed systems up to a system-size-dependent timescale. Interestingly, within such time scales, $\mathcal{D}(t)$ remains nearly independent of $\gamma$. Therefore, the $\gamma$ dependence of the $\mathcal{D}_{\text{dc}}$ can be seen as a potential signature of the inevitable finite-size effects at long times in open system techniques. We attribute the limitation of boundary-driven techniques to the improper order of limits inherent within this method. Specifically, in the boundary-driven technique, by construction, we take the limit in the order $\lim_{L\to\infty}\lim_{t\to\infty}$ instead of the physically appropriate order $\lim_{t\to\infty}\lim_{L\to\infty}$. The non-commutativity of these two order of limits inevitably leads to an onset of unavoidable finite-size effects that contaminate the NESS results. Based on our numerical results, we also present a somewhat intuitive argument why finite-size effects are inevitable in boundary-driven techniques.
The rest of the paper is organized as follows. In Sec. II, we introduce the models for which we perform our numerical analysis. In Sec. III, we present a comprehensive discussion of some of the key concepts related to this paper, namely the basic ideas of boundary-driven techniques and the extraction of diffusion constants from both, the boundary-driven open-system based method and from closed-system dynamics. In Sec. IV, we discuss the numerical methods we use for this work. This is followed by Sec. V where we present our main numerical results and, finally, we present our conclusion and future visions in Sec. VI.
@@
where $S^{x,y,z}_{r}$ are spin-$1/2$ operators at site $r$, $J>0$ is the antiferromagnetic coupling constant and $\Delta$ denotes the anisotropy in $z$ direction. We note that the sign of $J$ for this work does not really matter since we shall be focusing on infinite-temperature physics. The total number of sites is denoted by $L$ and throughout our work we focus on open boundary conditions. In this system the global magnetization is conserved, i.e. $[H,\sum_{r}S_{r}^{z}]=0$. Note that this global conservation of the total spin essentially makes the study of spin transport a meaningful and well-defined question. It is well-known that $H$ is integrable in terms of the Bethe Ansatz for any value of $\Delta$ and exhibits rich transport properties at high temperatures ranging from ballistic ($\Delta<1$) over superdiffusive ($\Delta=1$) to diffusive ($\Delta>1$) transport behavior [5, 39, 40]. For our further investigations, we focus on $\Delta\in\{1.5,3.0\}$ corresponding to the easy-axis regime where spin transport in the integrable system becomes diffusive. Later, we also consider a second system $H^{\prime}$ with
@@
where the additional next-to-nearest neighbor interactions of strength $\Delta^{\prime}>0$ breaks the integrability of the system and the situation becomes more generic.
@@ III.1 The boundary-driven system: Set-up and basics
[…]
$L_{1}=S^{+}_{1},\qquad$ $\alpha_{1}=\gamma(1+\mu),$ (5)
@@
$L_{3}=S^{+}_{L},\qquad$ $\alpha_{3}=\gamma(1-\mu),$
@@
Figure 2: Time evolution in the open system for the integrable spin-1/2 XXZ chain with anisotropy $\Delta=1.5$ and $L=20$ sites, as resulting for strong system-bath coupling $\gamma/J=1.0$ and weak driving $\mu=0.1$. Numerical data are obtained from exact stochastic unraveling (SU) with $N_{\text{traj}}=50000$ trajectories and time-evolving block decimation (TEBD) with bond dimension $\chi=150$. (a) Local densities, (b) local currents, and (c) steady-state profile at time $tJ=200$. In (b) average over times $tJ=100\dots 200$ for the shown bulk sites of the SU data is indicated by the yellow line. In (c) the linear function $f(r)$ results from fitting over bulk sites $r=5..15$ of SU data at times $tJ=200$.
It is straightforward to see from Eq. (5) that the Lindblad jump operators model baths which, in the absence of Hamiltonian, i.e. for the case of single-site driving, induce a net magnetization $\mu$ at the left edge (first site) and $-\mu$ at the right edge (last site) [5]. Such a set-up leads to a NESS with position-dependent magnetization along the chain and a non-zero time-independent current $j_{\text{ness}}$ for a finite system. Note that, for $\mu=0$, one has a trivial steady state $\rho\sim\mathbb{1}$, i.e. an infinite-temperature state, and one can interpret Eq. (5) as spin driving at infinite temperature. For non-zero $\mu$, the NESS current is finite and is the main observable [5]. Also, note that the linear response regime, for such a set-up, demands that $\mu$ should be chosen such that the NESS current $j_{\text{ness}}/\mu$ is independent of $\mu$, leading to desirable values of $\mu\ll 1$.
@@ III.2 Extracting diffusion constants from boundary-driven systems
The expression for the current in Eq. (7) can be derived from the lattice version of the continuity equation for the global conservation of the total spin i.e. , $[H,\sum_{r}S_{r}^{z}]=0$, as discussed in [5].
@@
with $t>0$. Note that, at $t=0$, $\mathcal{D}_{r}(t)$ is not defined, due to $\langle j_{r}(0)\rangle=0$ and $\langle\nabla S_{r}^{z}(0)\rangle=0$. However, as we will see later, $\lim_{t\to 0}\mathcal{D}_{r}(t)=0$. The quantity $\mathcal{D}_{r}(t)$ plays a central role in our work and has attracted less attention so far.
@@
Figure 4: Open-system diffusion constant $\mathcal{D}_{\text{dc}}$ versus system-bath coupling strength $\gamma$ for the integrable spin-1/2 XXZ chain with anisotropy (a) $\Delta=1.5$ and (b) $\Delta=3.0$, as resulting for weak driving $\mu=0.1$. Analogously, in (c) diffusion constants for the nonintegrable system with $\Delta^{\prime}=0.5$ and $\Delta=1.5$ are depicted. Numerical data from exact stochastic unraveling (SU) are shown for two system sizes $L=10,20$. Numerical data from TEBD calculations up to system sizes $L=50$ are also indicated. For comparison, numerical data for the closed-system diffusion constant according to the Kubo formula are depicted in (a) [41], (b) [42], and (c) [43].
@@ III.3 Extracting diffusion constants from closed systems
[…]
@@ IV Numerical Methods
As already mentioned earlier, solving the GKSL master equation numerically is a formidable task since the complexity grows as $d^{2L}$, where $d$ denotes the local Hilbert-space dimension. To this, we resort to two very powerful state-of-the-art numerical techniques to find well-controlled approximated solution of the GKSL master equations as discussed below.
@@ IV.1 Stochastic unraveling
The method of stochastic unraveling (SU), often also dubbed as Monte Carlo wavefunction technique, is powerful to get a good approximate solution of the GKSL master equation. Like any Monte Carlo method, here also the results are exact in the limit $N_{\text{sample}}\rightarrow\infty$ and we will shortly see the meaning of ‘sample’ for this method.
Stochastic unraveling essentially aims at propagating pure quantum states $|\psi\rangle$ instead of density matrices [35, 45, 46, 47], leading to different trajectories. In the end, the full density matrix is constructed by the average over these individual trajectories forming a stochastic solution of the GKSL master equation. The approach consists of an alternating scheme: A deterministic time evolution generated by an effective non-Hermitian Hamiltonian $H_{\mathrm{{eff}}}$, interspersed with stochastic quantum jumps with one of the Lindblad operators at specific times, which now will be discussed in more detail.
@@
$|\Psi(0)\rangle\propto\sum_{n=1}^{\text{dim}}c_{j}|j\rangle,$ (19)
@@ IV.2 Matrix product state (MPS) based technique
As a powerful alternative to stochastic unraveling, we resort to matrix product state (MPS) based technique to solve the GKSL master equation. In particular, we resort to time-evolving block decimation method (TEBD) [48] to time evolve the vectorized density matrix $\rho_{\text{vec}}(t)$ written as a MPS [29, 5, 49, 50]. More precisely, the GKSL master Eq. (3) can be recast in the vectorized form as
@@
where $\mathcal{L}_{\text{sup}}$ stands for the superoperator representation of the Liouvillian having the form $\mathcal{L}_{\text{sup}}=-i(H\otimes\mathbb{1}-\mathbb{1}\otimes H^{T})+\sum_{j}\frac{\alpha{j}}{2}(2L_{j}\otimes L_{j}^{*}-L_{j}^{\dagger}L_{j}\otimes\mathbb{1}-\mathbb{1}\otimes L_{j}^{*}L_{j}^{T})$ , where superscript $T$ and * denotes transpose and complex conjugation respectively. Note that $\mathbb{1}$ is the identity matrix with dimension dim. For our TEBD calculation, we use bond dimension $\chi=150$ and time step $dt=0.2$, we have carefully checked that all our results are well converged both w.r.t bond dimension and time step. The TEBD is implemented using the ITensor (Julia) library [51].

### v2

# Boundary-driven magnetization transport in the spin-$1/2$ XXZ chain: Role of the system-bath coupling strength and timescales

@@ Abstract
Understanding the transport properties of quantum many-body systems is a central challenge in condensed matter and statistical physics. Theoretical studies usually rely on two main approaches: Dynamics of linear-response functions in closed systems and boundary-driven dynamics governed by Markovian master equations for open systems. While the equivalence of their dynamical behavior has been explored in recent studies, a systematic comparison of the transport coefficients obtained from these two classes of approaches remains a largely open question. Here, we address this gap by comparing and contrasting the dc diffusion constant $\mathcal{D}_{\text{dc}}$ according to the two approaches, focusing on the specific example of magnetization transport in the spin-$1/2$ XXZ chain. Using exact numerical simulations for finite system sizes, we find (i) a clear mismatch between the two $\mathcal{D}_{\text{dc}}$ and (ii) a strong dependence of $\mathcal{D}_{\text{dc}}$ on the system-bath coupling strength for the open system, where neither (i) nor (ii) tend to vanish in the thermodynamic limit. These findings suggest limitations of the open-system approach to transport coefficients. To gain insight into the origin of (i) and (ii), we go beyond $\mathcal{D}_{\text{dc}}$ and analyze the full time dependence of the diffusion coefficient $D(t)$ in the open system. In this way, we find that both (i) and (ii) vanish up to a finite time scale. While this time scale gradually increases with system size and tends to be macroscopic in the thermodynamic limit, this increase is still slow compared to the increase of the time to reach the steady state, where (i) and (ii) do not vanish. This observation can be seen as a wrong, yet unavoidable order of limits of long times first and large system sizes afterwards.
@@ I Introduction
Over the last few decades, studying nonequilibrium dynamics in quantum many-body systems has become a key area of research in both theoretical and experimental condensed matter physics, with broad ramifications ranging from fundamental questions in statistical mechanics to advanced applications in quantum control and quantum computation [1, 2, 3, 4]. One of the most ubiquitous manifestations of nonequilibrium behavior are transport processes of conserved quantities such as energy, particle number and magnetization, which are relevant for both closed/isolated and open quantum systems coupled to external environments. A comprehensive understanding of transport is thus crucial for elucidating the emergent macroscopic behavior as well as the underlying microscopic mechanisms governing quantum many-body dynamics such as the build up of nonequilibrium steady states (NESS) and the relaxation to them [5]. Over the past years, notable progress in this field has been driven by the development of new theoretical concepts like eigenstate thermalization hypotheses [6, 7, 8, 9] and dynamical quantum typicality [10, 11, 12, 13, 14], as well as sophisticated analytical and numerical methods and advanced experimental techniques. We now provide a brief overview of some of the key numerical and analytical methods commonly employed to study such problems, which are particularly relevant to the present work.
@@
On the numerical frontier, an alternative to traditional LRT-based closed-system technique is given within the framework of boundary-driven open quantum systems. In such a setup, the system is coupled to an environment, for instance, modeled by external baths at different temperatures or chemical potentials [26, 27, 28, 29, 30]. This coupling drives the system out of equilibrium and induces transport, typically resulting in a NESS with a constant current in the bulk and a characteristic density profile in the long-time limit. A widely employed approach for describing the dynamics of such open-system scenarios is provided by Gorini–Kossakowski–Sudarshan–Lindblad (GKSL) quantum master equation [31, 32]. While the derivation of the GKSL master equation from a microscopic system-bath model is a nontrivial task [33, 34], it provides the most general form of a time-local (Markovian) master equation whose dynamical map is completely positive and trace-preserving (CPTP), thus ensuring the evolution of any physical density matrix into another valid density matrix [35]. This formalism offers considerable advantages for numerical simulations, particularly in conjunction with stochastic methods or matrix-product-state techniques based upon vectorized density matrices.
@@
Figure 1: Sketch summarizing our key findings for the time-dependent diffusion constant $\mathcal{D}(t)$ from the open-system approach (purple). The closed-system Kubo-based approach [41, 42, 43] is indicated (orange). An early-time plateau shows a (nearly) $\gamma$-independent agreement up to a system-size–dependent scale $t_{\mathrm{fs}}J\propto L$. At long times, the NESS-based dc values disagree with Kubo and strongly depend on the system–bath coupling $\gamma$.
Using exact numerical simulations for finite system sizes, we find (i) a clear mismatch between the two $\mathcal{D}_{\text{dc}}$ and (ii) a strong dependence of $\mathcal{D}_{\text{dc}}$ on the system-bath coupling strength for the open system, where neither (i) nor (ii) tend to vanish in the thermodynamic limit. These findings suggest limitations of the open-system approach to transport coefficients. To gain insight into the origin of (i) and (ii), we go beyond $\mathcal{D}_{\text{dc}}$ and analyze the full time dependence of the diffusion coefficient $D(t)$ in the open system, as sketched in Fig.1. In this way, we find that both (i) and (ii) vanish up to a finite time scale. While this time scale gradually increases with system size and tends to be macroscopic in the thermodynamic limit, this increase is still slow compared to the increase of the time to reach the steady state, where (i) and (ii) do not vanish. This observation can be seen as a wrong, yet unavoidable order of limits of long times first and large system sizes afterwards.
The rest of the paper is organized as follows. In Sec. II, we introduce the models for which we perform our numerical analysis. In Sec. III, we present a comprehensive discussion of some of the key concepts related to this paper, namely the basic ideas of boundary-driven techniques and the extraction of diffusion constants from both, the boundary-driven open-system based method and from closed-system dynamics. In Sec. IV, we discuss the numerical methods we use for this work. This is followed by Sec. V where we present our main numerical results and, finally, we present our conclusion in Sec. VI.
@@
where $S^{x,y,z}_{r}$ are spin-$1/2$ operators at site $r$, $J>0$ is the antiferromagnetic coupling constant and $\Delta$ denotes the anisotropy in $z$ direction. We note that the sign of $J$ for this work does not really matter since we shall be focusing on infinite-temperature physics. The total number of sites is denoted by $L$ and throughout our work we focus on open boundary conditions. In this system the global magnetization is conserved, i.e. $H,\sum_{r}S_{r}^{z}]=0$. Note that this global conservation of the total spin essentially makes the study of spin transport a meaningful and well-defined question. It is well-known that $H$ is integrable in terms of the Bethe Ansatz for any value of $\Delta$ and exhibits rich transport properties at high temperatures ranging from ballistic ($\Delta<1$) through superdiffusive ($\Delta=1$) to diffusive ($\Delta>1$) transport behavior [5, 39, 40]. For our further investigations, we focus on $\Delta\in\{1.5,3.0\}$, corresponding to the easy-axis regime where spin transport in the integrable system becomes diffusive. Furthermore, we also consider a second system $H^{\prime}$ with
@@
where the additional next-to-nearest neighbor interactions of strength $\Delta^{\prime}>0$ breaks the integrability of the system and the model becomes more generic.
@@ III.1 The boundary-driven system: Set-up and basics
where $\{\cdot,\cdot\}$ is the anticommutator. Moreover, $\alpha_{j}$ and $L_{j}$ denote non-negative rates and the Lindblad jump operators, respectively. Here, we consider the jump operators of the following form [5]:
$L_{1}=S^{+}_{1}$$\qquad$ $\alpha_{1}=\gamma(1+\mu),$ (5)
@@
$L_{3}=S^{+}_{L}$$\qquad$ $\alpha_{3}=\gamma(1-\mu),$
@@
Figure 3: Time evolution in the open system for the integrable spin-1/2 XXZ chain with anisotropy $\Delta=1.5$ and $L=20$ sites, as resulting for strong system-bath coupling $\gamma/J=1.0$ and weak driving $\mu=0.1$. Numerical data are obtained from exact stochastic unraveling (SU) with $N_{\text{traj}}=50000$ trajectories and time-evolving block decimation (TEBD) with bond dimension $\chi=150$. (a) Local densities, (b) local currents, and (c) steady-state profile at time $tJ=200$. In (b) the average over the time window $tJ=100\dots 200$ for the shown bulk sites of the SU data is indicated by the yellow line. Panel (c) depicts the linear function $f(r)$ from fitting over bulk sites $r=5..15$ of SU data at times $tJ=200$.
It is straightforward to see from Eq. (5) that the Lindblad jump operators model baths which, in the absence of Hamiltonian, i.e. for the case of single-site driving, induce a net magnetization $\mu$ at the left edge (first site) and $\mu$ at the right edge (last site) [5]. Such a set-up leads to a NESS with position-dependent magnetization along the chain and a non-zero time-independent current $j_{\text{ness}}$ for a finite system. Note that, for $\mu=0$, one has a trivial steady state $\rho\sim\mathbb{1}$, i.e. an infinite-temperature state, and one can interpret Eq. (5) as spin driving at infinite temperature. For non-zero $\mu$, the NESS current is finite and is the main observable [5]. Also, note that the linear response regime demands that $\mu$ should be chosen such that the NESS current $j_{\text{ness}}/\mu$ is independent of $\mu$, leading to desirable values of $\mu\ll 1$.
@@ III.2 Extracting diffusion constants from boundary-driven systems
The expression for the current in Eq. (7) can be derived from the lattice version of the continuity equation for the global conservation of the total spin i.e. , $H,\sum_{r}S_{r}^{z}]=0$, as discussed in [5].
@@
with $t>0$. Note that, at $t=0$, $\mathcal{D}_{r}(t)$ is not defined, due to $\langle j_{r}(0)\rangle=0$ and $\langle\nabla S_{r}^{z}(0)\rangle=0$. However, as we will see later, $\lim_{t\to 0}\mathcal{D}_{r}(t)=0$.
@@
Figure 5: Open-system diffusion constant $\mathcal{D}_{\text{dc}}$ versus system-bath coupling strength $\gamma$ for the integrable spin-1/2 XXZ chain with anisotropy (a) $\Delta=1.5$ and (b) $\Delta=3.0$, as resulting for weak driving $\mu=0.1$. Analogously, in (c) diffusion constants for the nonintegrable system with $\Delta^{\prime}=0.5$ and $\Delta=1.5$ are depicted. Numerical data from exact stochastic unraveling (SU) are shown for two system sizes $L=10,20$. Numerical data from TEBD calculations up to system sizes $L=50$ are also indicated. For comparison, numerical data for the closed-system diffusion constant according to the Kubo formula are depicted in (a) [41], (b) [42], and (c) [43]. Further, open-system diffusion constants $D_{r}(t)$ extracted at (representative) finite times $tJ=20$ from Figs. 6, 7 are indicated in purple.
@@ III.3 Extracting diffusion constants from closed systems
$\mathcal{D}_{\text{closed}}(t)=\frac{1}{\chi_{s}}\int_{0}^{t}dt^{\prime}\,\langle\mathcal{J}_{S}(t^{\prime})\mathcal{J}_{S}\rangle,$ (12)
@@ IV Numerical Methods
As already mentioned earlier, solving the GKSL master equation numerically is a formidable task since the complexity grows as $d^{2L}$, where $d$ denotes the local Hilbert-space dimension. We resort to two powerful state-of-the-art numerical techniques to find well-controlled approximated solution of the GKSL master equations as discussed below.
@@ IV.1 Stochastic unraveling
The method of stochastic unraveling (SU), often also dubbed as Monte Carlo wavefunction technique, is a powerful technique to get a good approximate solution of the GKSL master equation. Like any Monte Carlo method, here the results are exact in the limit $N_{\text{sample}}\rightarrow\infty$ and we will shortly see the meaning of ‘sample’ for this method.
Stochastic unraveling essentially aims at propagating pure quantum states $|\psi\rangle$ instead of density matrices [35, 45, 46, 47], leading to different trajectories. In the end, the full density matrix is constructed by the average over these individual trajectories forming a stochastic solution of the GKSL master equation. The approach consists of an alternating scheme: A deterministic time evolution generated by an effective non-Hermitian Hamiltonian $H_{\mathrm{{eff}}}$, interspersed with stochastic quantum jumps with one of the Lindblad operators at specific times.
@@
$|\Psi(0)\rangle\propto\sum_{n=1}^{\text{dim}(H)}c_{j}|j\rangle,\$(19)
@@ IV.2 Matrix product state (MPS) based technique
As an alternative to stochastic unraveling, we resort to matrix product state (MPS) based technique to solve the GKSL master equation. In particular, we resort to time-evolving block decimation method (TEBD) [48] to time evolve the vectorized density matrix $\rho_{\text{vec}}(t)$ written as a MPS [29, 5, 49, 50]. More precisely, the GKSL master Eq. (3) can be recast in the vectorized form as
@@
where $\mathcal{L}_{\text{sup}}$ stands for the superoperator representation of the Liouvillian having the form $\mathcal{L}_{\text{sup}}=-i(H\otimes\mathbb{1}-\mathbb{1}\otimes H^{T})+\sum_{j}\frac{\alpha{j}}{2}(2L_{j}\otimes L_{j}^{*}-L_{j}^{\dagger}L_{j}\otimes\mathbb{1}-\mathbb{1}\otimes L_{j}^{*}L_{j}^{T})$ , where superscript $T$ and * denotes transpose and complex conjugation respectively. Note that $\mathbb{1}$ is the identity matrix with dimension $\text{dim}(H)$. For our TEBD calculation, we use bond dimension $\chi=150$ and time step $dt=0.2$, we have carefully checked that all our results are well converged both w.r.t bond dimension and time step. The TEBD is implemented using the ITensor (Julia) library [51].